# Zero-Shot Image Restoration Using Few-Step Guidance of Consistency Models (and Beyond)

### v1

@@ Abstract
In recent years, it has become popular to tackle image restoration tasks with a single pretrained diffusion model (DM) and data-fidelity guidance, instead of training a dedicated deep neural network per task. However, such “zero-shot” restoration schemes currently require many Neural Function Evaluations (NFEs) for performing well, which may be attributed to the many NFEs needed in the original generative functionality of the DMs. Recently, faster variants of DMs have been explored for image generation. These include Consistency Models (CMs), which can generate samples via a couple of NFEs. However, existing works that use guided CMs for restoration still require tens of NFEs or fine-tuning of the model per task that leads to performance drop if the assumptions during the fine-tuning are not accurate. In this paper, we propose a zero-shot restoration scheme that uses CMs and operates well with as little as 4 NFEs. It is based on a wise combination of several ingredients: better initialization, back-projection guidance, and above all a novel noise injection mechanism. We demonstrate the advantages of our approach for image super-resolution, deblurring and inpainting. Interestingly, we show that the usefulness of our noise injection technique goes beyond CMs: it can also mitigate the performance degradation of existing guided DM methods when reducing their NFE count.
Figure 1: Super-resolution ×$\times 4$ with bicubic kernel and noise level of 0.05. From left to right and top to bottom: original, observation, DiffPIR [39] (20 NFEs), DDRM [15] (20 NFEs), and our CM4IR (4 NFEs).
Figure 2: Deblurring with Gaussian kernel and noise level of 0.025. From left to right and top to bottom: original, observation, DiffPIR [39] (20 NFEs), DDRM [15] (20 NFEs), and our CM4IR (4 NFEs).
@@ 1 Introduction
Since the breakthroughs in deep learning a decade ago [16], it has become common to address these problems by training a different deep neural network (DNN) for each predefined observation model in a supervised manner [7, 17]. Specifically, given a collection of high-quality images $\{\mathbf{x}^{*}_{i}\}$, training pairs $\{\mathbf{y}_{i},\mathbf{x}^{*}_{i}\}$ are generated by mapping $\mathbf{x}^{*}_{i}$ to $\mathbf{y}_{i}$ using (1), and a DNN is trained to invert the map. However, these “task-specific” DNNs suffer from a huge performance drop when the observations at test-time mismatch (even slightly) the assumptions made in training [21, 29, 2], which limits their applicability in many practical cases.
@@ 2.2 Restoration via Guidance of DMs
which is used, e.g., in [6]. Yet, acceleration (less NFEs) is obtained with the back-projection (BP) guidance [28, 30, 31]
@@
where $\mathbf{A}^{\dagger}=\mathbf{A}^{T}(\mathbf{A}\mathbf{A}^{T})^{-1}$ is the pseudoinverse of $\mathbf{A}$, which can be regularized and oftentimes has efficient implementation (generally by conjugate gradients [11], and in super-resolution and deblurring even by FFT [9]). This guidance has been used recently in [25, 23, 33, 9], sometimes under different names (e.g., “pseudoinverse guidance”). When the SVD of $\mathbf{A}$ can be computed efficiently, a guidance that resembles BP but can mitigate noise amplification per singular component has been used in the DDRM method [15]. Importantly, existing zero-shot restoration techniques based on guided DMs still require at least a couple of dozens NFEs.
@@ 3 The Proposed Method
Table 1: Ablation study on super-resolution. PSNR [dB] ($\uparrow$) and LPIPS ($\downarrow$) results on LSUN Bedroom validation set.
@@
Table 2: Super-resolution and inpainting. PSNR [dB] ($\uparrow$) and LPIPS ($\downarrow$) results on LSUN Bedroom validation set.
@@
Gauss. Deblurring $\sigma_{y}$=0.025 26.07 / 0.339 19.74 / 0.342 28.94 / 0.221 27.48 / 0.319 28.85 / 0.217
Gauss. Deblurring $\sigma_{y}$=0.05 24.18 / 0.453 19.08 / 0.543 27.35 / 0.280 26.14 / 0.363 27.37 / 0.270
@@
Gauss. Deblurring $\sigma_{y}$=0.025 26.69 / 0.346 N/A 29.84 / 0.258 27.93 / 0.330 29.62 / 0.246
Gauss. Deblurring $\sigma_{y}$=0.05 24.54 / 0.453 N/A 28.33 / 0.316 26.64 / 0.370 27.76 / 0.295
@@ 4 Experiments
Datasets. We consider the datasets LSUN bedroom and LSUN cat (256$\times$256), for which there exits both well pretrained CMs [27] and DMs (specifically DDPMs) [12]. For the test images, we consider 300 validation images for each dataset (as done in [38]).
@@ 4.2 Comparison with other methods
We compare our CM4IR also with zero-shot guided DM approaches: DDRM [15] and DiffPIR [39], which require 20 NFEs. We apply these methods with the same DMs.
@@ 5 Conclusion
Interestingly, we also showed that the usefulness of our noise injection technique is not limited to CMs. Specifically, we applied it to improve the performance of existing guided DM methods when drastically reducing NFE number. Yet, these variants did not outperform our CM4IR. Another interesting take-away observation from our study is the limitations of using a task-adapted technique (as used in CoSIGN), which does not seem to generalize well to conditions at test time that mismatch the configuration that has been used during the fine-tuning/adaptation. Alleviating this issue is an interesting direction for future research.
@@ Acknowledgment
The work is supported by ISF grant no. 1940/23.
@@ B.1 Hyperparameter setting
Bicub. SRx4 $\sigma_{e}$=0.05 $\eta=0.1$, $i_{N}=250$, $\gamma=0.2$, $\delta_{i}=(0.0,0.3,0.1,0.0)$ $\eta=0.1$, $i_{N}=250$, $\gamma=0.2$, $\delta_{i}=(0.1,0.1,0.0,0.0)$
Gauss. Deblurring $\sigma_{e}$=0.025 $\eta=0.1$, $i_{N}=75$, $\gamma=0.01$, $\zeta=3.0$, $\delta_{i}=(0.1,0.0,0.0,0.05)$ $\eta=0.1$, $i_{N}=125$, $\gamma=0.05$, $\zeta=4.0$, $\delta_{i}=(0.0,0.0,0.0,0.0)$
Gauss. Deblurring $\sigma_{e}$=0.05 $\eta=0.1$, $i_{N}=150$, $\gamma=0.05$, $\zeta=1.0$, $\delta_{i}=(0.1,0.0,0.0,0.05)$ $\eta=0.1$, $i_{N}=250$, $\gamma=0.2$, $\zeta=1.0$, $\delta_{i}=(0.0,0.0,0.0,0.0)$
@@
B.2 Effectiveness of $\delta$
@@
Gauss. Deblurring $\sigma_{y}$=0.025 28.80 / 0.219 28.85 / 0.217
Gauss. Deblurring $\sigma_{y}$=0.05 27.33 / 0.271 27.37 / 0.270
@@
B.3 Different amount of NFEs
@@
Task NFEs, $\tau_{n}$} 3 NFEs 4 NFEs 5 NFEs
@@
B.4 Advantages over additional competitors
@@
B.5 Reducing NFEs of additional DM-based method
@@
B.6 Additional inpainting results
@@
B.7 More qualitative results

### v2

@@ Abstract
In recent years, it has become popular to tackle image restoration tasks with a single pretrained diffusion model (DM) and data-fidelity guidance, instead of training a dedicated deep neural network per task. However, such “zero-shot” restoration schemes currently require many Neural Function Evaluations (NFEs) for performing well, which may be attributed to the many NFEs needed in the original generative functionality of the DMs. Recently, faster variants of DMs have been explored for image generation. These include Consistency Models (CMs), which can generate samples via a couple of NFEs. However, existing works that use guided CMs for restoration still require tens of NFEs or fine-tuning of the model per task that leads to performance drop if the assumptions during the fine-tuning are not accurate. In this paper, we propose a zero-shot restoration scheme that uses CMs and operates well with as little as 4 NFEs. It is based on a wise combination of several ingredients: better initialization, back-projection guidance, and above all a novel noise injection mechanism. We demonstrate the advantages of our approach for image super-resolution, deblurring and inpainting. Interestingly, we show that the usefulness of our noise injection technique goes beyond CMs: it can also mitigate the performance degradation of existing guided DM methods when reducing their NFE count. The source code is available at https://github.com/tirer-lab/CM4IR
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 1: Super-resolution ×$\times 4$ with bicubic kernel and noise level of 0.05. From left to right and top to bottom: original, observation, DPS [6] (1000 NFEs), DiffPIR [40] (20 NFEs), DDRM [15] (20 NFEs) and our CM4IR (4 NFEs).
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 2: Deblurring with Gaussian kernel and noise level of 0.025. From left to right and top to bottom: original, observation, DPS [6] (1000 NFEs), DiffPIR [40] (20 NFEs), DDRM [15] (20 NFEs) and our CM4IR (4 NFEs).
@@ 1 Introduction
Since the breakthroughs in deep learning a decade ago [16], it has become common to address these problems by training a different deep neural network (DNN) for each predefined observation model in a supervised manner [7, 17]. Specifically, given a collection of high-quality images $\{\mathbf{x}^{*}_{i}\}$, training pairs $\{\mathbf{y}_{i},\mathbf{x}^{*}_{i}\}$ are generated by mapping $\mathbf{x}^{*}_{i}$ to $\mathbf{y}_{i}$ using (1), and a DNN is trained to invert the map. However, these “task-specific” DNNs suffer from a huge performance drop when the observations at test-time mismatch (even slightly) the assumptions made in training [21, 29, 2, 32], which limits their applicability in many practical cases.
@@ 2.2 Restoration via Guidance of DMs
which is used, e.g., in [6]. Yet, acceleration (less NFEs) is obtained with the back-projection (BP) guidance [28, 30]
@@
where $\mathbf{A}^{\dagger}=\mathbf{A}^{T}(\mathbf{A}\mathbf{A}^{T})^{-1}$ is the pseudoinverse of $\mathbf{A}$, which can be regularized and oftentimes has efficient implementation (generally by conjugate gradients [11], and in super-resolution and deblurring even by FFT [9]). See [31] for theoretical analysis of the acceleration property. This guidance has been used recently in [25, 23, 34, 9], sometimes under different names (e.g., “pseudoinverse guidance”). When the SVD of $\mathbf{A}$ can be computed efficiently, a guidance that resembles BP but can mitigate noise amplification per singular component has been used in the DDRM method [15]. Importantly, existing zero-shot restoration techniques based on guided DMs still require at least a couple of dozens NFEs.
@@ 3 The Proposed Method
Table 1: Ablation study on super-resolution with 4 NFEs. PSNR [dB] ($\uparrow$) and LPIPS ($\downarrow$) results on LSUN Bedroom validation set.
@@
Table 2: Super-resolution, deblurring and inpainting. PSNR [dB] ($\uparrow$) and LPIPS ($\downarrow$) results on LSUN Bedroom validation set.
@@
Gauss. Deblurring $\sigma_{y}$=0.025 26.07 / 0.339 19.74 / 0.342 28.94 / 0.221 27.48 / 0.319 29.00 / 0.213
Gauss. Deblurring $\sigma_{y}$=0.05 24.18 / 0.453 19.08 / 0.543 27.35 / 0.280 26.14 / 0.363 27.42 / 0.258
@@
Table 3: Super-resolution, deblurring and inpainting. PSNR [dB] ($\uparrow$) and LPIPS ($\downarrow$) results on LSUN Cat validation set.
@@
Gauss. Deblurring $\sigma_{y}$=0.025 26.69 / 0.346 N/A 29.84 / 0.258 27.93 / 0.330 29.76 / 0.247
Gauss. Deblurring $\sigma_{y}$=0.05 24.54 / 0.453 N/A 28.33 / 0.316 26.64 / 0.370 28.06 / 0.299
@@ 4 Experiments
Datasets. We consider the datasets LSUN bedroom and LSUN cat (256$\times$256), for which there exits both well pretrained CMs [27] and DMs (specifically DDPMs) [12]. For the test images, we consider 300 validation images for each dataset (as done in [39]). In the appendix we also report results for ImageNet 64$\times$64.
@@ 4.2 Comparison with other methods
We compare our CM4IR also with zero-shot guided DM approaches: DDRM [15] and DiffPIR [40], which require 20 NFEs. We apply these methods with the same DMs downloaded from https://huggingface.co/google.
@@ 5 Conclusion
Interestingly, we also showed that the usefulness of our noise injection technique is not limited to CMs. Specifically, we applied it to improve the performance of existing guided DM methods when drastically reducing NFE number. Yet, these variants did not outperform our CM4IR.
@@ Acknowledgment
The work is supported by ISF (No. 1940/23) and MOST (No. 0007091) grants.
@@ B.1 Hyperparameter setting
Bicub. SRx4 $\sigma_{e}$=0.05 $\eta=0.1$, $i_{N}=250$, $\gamma=0.2$, $\delta_{i}=(0.0,0.3,0.05,0.1)$ $\eta=0.1$, $i_{N}=250$, $\gamma=0.2$, $\delta_{i}=(0.1,0.1,0.0,0.0)$
Gauss. Deblurring $\sigma_{e}$=0.025 $\eta=0.1$, $i_{N}=90$, $\gamma=0.02$, $\zeta=3.0$, $\delta_{i}=(0.0,0.0,0.0,0.0)$ $\eta=0.1$, $i_{N}=100$, $\gamma=0.03$, $\zeta=4.0$, $\delta_{i}=(0.0,0.0,0.0,0.0)$
Gauss. Deblurring $\sigma_{e}$=0.05 $\eta=0.1$, $i_{N}=160$, $\gamma=0.07$, $\zeta=1.5$, $\delta_{i}=(0.0,0.0,0.0,0.0)$ $\eta=0.1$, $i_{N}=180$, $\gamma=0.1$, $\zeta=2.0$, $\delta_{i}=(0.0,0.0,0.0,0.0)$
@@
B.2 Robustness to hyperparameter settings
To asses the robustness of CM4IR to the hyperparameter settings, we examine PSNR/LPIPS for SRx4 with $\sigma_{y}$=0.025 on the LSUN bedroom when we moderately modify the parameters from the values stated in B.1. Figure 8 (left) shows that increasing $\eta$ has only minor effect on the PSNR/LPIPS. Consistent PSNR/LPIPS are also observed in Figure 8 (right) when decreasing $(i_{N},\gamma)$. Recall that $i_{N}$ determines the first noise level (higher $i_{N}$ implies stronger noise) and $\gamma$ determines its decay rate. Thus, it makes sense to pair them, as examined.
Figure 8: PSNR/LPIPS for different parameters. The chosen values are in bold.
B.3 Effectiveness of $\delta$
@@
B.4 Different amount of NFEs
@@
Task NFEs 3 NFEs 4 NFEs 5 NFEs
@@
B.5 Advantages over additional competitors
@@
B.6 Reducing NFEs of additional DM-based method
@@
B.7 Additional inpainting results
@@
Original
Observation
DiffPIR
DDRM
CM (40)
CoSIGN
CM4IR (ours)
@@
B.8 Additional dataset
In this section, we evaluate CM4IR also for ImageNet 64$\times$64 for SR with bicubic downsampling of factor 2 and different noise levels. This factor allows to obtain meaningful recoveries despite the low resolution of the original 64$\times$64 images. The results are presented in Table 11, and show that CM4IR outperforms the other methods in terms of PSNR and has competitive LPIPS despite using only 4 NFEs.
Table 11: Bicubic SRx2 for ImageNet 64x64. PSNR [dB] and LPIPS.
Task Method CM (40 NFEs) DDRM (20 NFEs) DiffPIR (20 NFEs) CM4IR (Ours, 4 NFEs)
SRx2, $\sigma_{y}$=0.01 27.33 / 0.165 29.26 / 0.104 28.64 / 0.132 30.11 / 0.122
SRx2, $\sigma_{y}$=0.025 25.98 / 0.235 28.38 / 0.134 27.10 / 0.163 29.20 / 0.148
SRx2, $\sigma_{y}$=0.05 24.51 / 0.309 27.02 / 0.178 23.65 / 0.250 27.74 / 0.188
B.9 More qualitative results